OpenStudy (anonymous):

rationalize the denominator 3√3y^4/3√6x^4

OpenStudy (anonymous):

\[\frac{3\sqrt{3y^4}}{3\sqrt{3x^4}} = \frac{y^2\sqrt{3}}{\sqrt{3x^4}}\]

OpenStudy (anonymous):

Or is it Cubed root?

OpenStudy (anonymous):